(x)^1/2-(x+3)^1/2=0
Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12844)

(x)^1/2-(x+3)^1/2=0
(x)^1/2=(x+3)^1/2
sqrtx=sqrt(x+3)
x=x+3
x-x=3
0=3 no unique solution for x.

Answer by solver91311(24713)






If you first add to both sides of the equation and re-write the fractional exponents as radicals:



You can readily see that this equation has no solution. That's because over the most restrictive domain (the left hand side function), namely

,

The values of the two sides of the equation will differ by a factor of at the low end of the domain and will differ by a decreasing amount as x gets larger, but the difference can never be exactly zero.

This is the reason that using ordinary solution methods on this equation results in an absurdity:



Square both sides:



Add -x to both sides:

Absurd.

Therefore, the solution set of the given equation is the empty set.

John